# Subtree Mode and Applications

### v1

@@ Abstract
The mode of a collection of values (i.e., the most frequent value in the collection) is a key summary statistic. Finding the mode in a given range of an array of values is thus of great importance, and constructing a data structure to solve this problem is in fact the well-known Range Mode problem. In this work, we introduce the Subtree Mode (SM) problem, the analogous problem in a leaf-colored tree, where the task is to compute the most frequent color in the leaves of the subtree of a given node. SM is motivated by several applications in domains such as text analytics and biology, where the data are hierarchical and can thus be represented as a (leaf-colored) tree. Our central contribution is a time-optimal algorithm for SM that computes the answer for every node of an input $N$-node tree in $\mathcal{O}(N)$ time. We further show how our solution can be adapted for node-colored trees, or for computing the $k$ most frequent colors, in the optimal $\mathcal{O}(N)$ time, for any given $k=\mathcal{O}(1)$. Moreover, we prove that a similarly fast solution for when the input is a sink-colored directed acyclic graph instead of a leaf-colored tree is highly unlikely. Our experiments on real datasets with trees of up to $7.3$ billion nodes demonstrate that our algorithm is faster than baselines by at least one order of magnitude and much more space efficient. Last, we present case studies showing the effectiveness of our approach in pattern mining and sequence-to-database search applications.
@@ 1 Introduction
A key summary statistic of a collection of values is its mode (i.e., the most frequent value in the collection) [HKP11]. Finding the mode in a given range of values of an array (e.g., a window of a sequence) is thus of great importance. In fact, constructing a data structure to solve this problem is the well-known Range Mode (RM) problem [KMS05, Cha+14, Gre+10, VX20]. For instance, RM allows finding the most frequently purchased item over a certain time period [HL23], the most frequent $q$-gram (i.e., length-$q$ substring) occurring in a genomic region [APP18], or the most frequent value of an attribute in a range of database tuples [HKP11].
@@
SM is motivated by several applications in domains such as text analytics and biology. We sketch some of these applications below.
SNP-based Phylogenetic Tree Annotation. Single-nucleotide polymorphisms (SNPs) are genetic variations at specific nucleotide sites within a species’ genome. SNPs are linked to diseases such as cancer, Alzheimer’s disease, and other inherited disorders [KLE11, TK04]. Owing to their physical proximity to disease-associated loci, SNP alleles are frequently co-inherited with pathogenic variants across generations, reflecting the principle of genetic linkage [Syv01]. Phylogenetic trees show evolutionary relationships between organisms; a leaf represents an organism from which a DNA sequence has been obtained, and an internal node represents a common ancestor of all the organisms that correspond to its leaf descendants [WL11]. The leaves are often annotated manually with categorical values related to SNPs (e.g., types of diseases [Liu+09]), and each internal node with a value summarizing the values of its subtree [WHN22, Sch+18, Pea+25, LB24]. We can model such a phylogenetic tree as $\mathcal{T}$ and color each leaf of $\mathcal{T}$ according to the SNP-related value of its corresponding leaf in the phylogenetic tree. Then, the mode of $v$ identifies the most prevalent SNP-related value among the group of organisms corresponding to $v$, supporting interpretation and hypothesis generation about evolutionary processes [Hua+25].
@@
Fig. 1 shows an SNP-based phylogenetic tree alike the ones used to study the Influenza A virus in [Liu+09]. The leaves are annotated with Influenza A subtypes (H1N1, H1N2, H5N1, and H5N8), and the internal nodes represent lineages (subtrees sharing specific patterns of SNPs). An internal node with .1 and .2 corresponds to a subtype in the Eastern and Western Hemisphere, respectively. We model the tree in Fig. 1 as $\mathcal{T}$ in our SM problem and assign colors 0 (blue), $1$ (green), $2$ (red), and $3$ (orange) to the leaves of subtype H1N1, H1N2, H5N1, and H5N8, respectively. Thus, the number of nodes $N$ of $\mathcal{T}$ is $20$, and the number of colors $\Delta$ is $4$. Given a query consisting of the node H1, SM outputs 0, as the subtree rooted at this node has more blue leaves than green.
Top-11 Document Retrieval ($1$-DR). In the $1$-DR problem [NN12, Hon+14, NN17, NN25], a collection $\mathcal{S}$ of $\Delta$ documents (strings), $S_{0},\ldots,S_{\Delta-1}$, is given for preprocessing, and we are asked to answer queries of the following type: given a query pattern $P$, output the string in $\mathcal{S}$ in which $P$ occurs most frequently as a substring. $1$-DR can be reduced to SM. We preprocess $\mathcal{S}$ by first constructing its suffix tree (i.e., the compacted trie of the suffixes of the string $S_{0}\mathdollar_{0}\ldots S_{\Delta-1}\mathdollar_{\Delta-1}$, where $\mathdollar_{i}$, for each $i\in[0,\Delta)$, is a unique delimiter) [Wei73] and then coloring the leaves corresponding to the suffixes starting in $S_{i}\mathdollar_{i}$ with color $i$. This is the leaf-colored tree in the constructed instance of SM. For a query pattern $P$, we spell $P$ on the suffix tree, arriving at a node $v$, compute the mode of $v$, say $i$, and output $S_{i}$ as the answer.
@@
Uniform Pattern Mining (UPM). Consider two strings, one comprised of male-targeted ads and another comprised of female-targeted ads, and that we mine a pattern “strong leader” which occurs much more frequently in the former string. Ads with this pattern may perpetuate gender stereotypes and influence how opportunities or messages are presented to different ad viewer groups, causing discrimination and bias in decision-making or perception. To prevent this, in the spirit of statistical parity [Dwo+12] [^1] This fairness measure requires the probability distributions of outcomes to be similar across all subpopulations of a population., we can mine patterns with “similar” frequencies in all strings of an input collection by formulating and solving the following problem, called Uniform Pattern Mining (UPM): Given a collection $\mathcal{S}$ of $\Delta$ strings, $S_{0},\ldots,S_{\Delta-1}$, and an integer $\varepsilon\geq 0$ specified based on domain knowledge, UPM asks for all strings (patterns) whose frequencies in any pair of strings in $\mathcal{S}$ differ by at most $\varepsilon$. When the strings in $\mathcal{S}$ represent different subpopulations of a user population and $\varepsilon$ is “small”, such patterns prevent the discrimination of these subpopulations. On the other hand, when $\varepsilon$ is “large”, those patterns with large differences in their frequencies reveal behavioral preferences that prevail in user subpopulations (e.g., they may represent movie genres viewed by much more men than women). The UPM problem is solved via a reduction to SM.
@@
Consistent Query String (CQS). In the CQS problem, a collection $\mathcal{S}$ of $\Delta$ strings, $S_{0},\ldots,S_{\Delta-1}$, is given for preprocessing, and we are asked to quantify how similar a query string $P$ is to the strings in $\mathcal{S}$. This can be achieved by counting the number of distinct $q$-grams $Q$ of $P$ whose frequency in $P$ is in the interval ${[\min_{S\in\mathcal{S}}|\textsf{occ}_{S}(Q)|-\varepsilon,}\max_{S\in\mathcal{S}}|\textsf{occ}_{S}(Q)|+\varepsilon]$, where $|\textsf{occ}_{S}(Q)|$ is the frequency of $Q$ in string $S$ and $\varepsilon$ is a user-specified parameter capturing approximate consistency. We call each such $q$-gram $\varepsilon$-consistent with $\mathcal{S}$. Clearly, $P$ is similar to the strings in $\mathcal{S}$ if most of its $q$-grams are $\varepsilon$-consistent with $\mathcal{S}$. The CQS problem is particularly relevant for databases of highly-similar strings, which are common in genomics, as they are constructed over collections with a shared evolutionary history or a common function [SM17, Mis+21]. The CQS problem can be solved via a reduction to SM. The values for $q$ and $\varepsilon$ can be set based on domain knowledge differently for each query.
@@
Consider two DNA sequences, one that is a genetic variant of SARS-Cov-2 [Nat24] and another that is a genetic variant of the Influenza A virus [Nat25]. Suppose that a biologist does not know whether each of these sequences is a genetic variant of SARS-CoV-2 and wants to check this. They can use the first sequence as query $P_{1}$ in a database $\mathcal{S}$ comprised of $2,000$ different genetic variants of the SARS CoV-2 virus [Nat24], and solve CQS for $\varepsilon=0$ and $q=4$. They will find that $99.9\%$ of the $q$-grams of $P_{1}$ are 0-consistent with $\mathcal{S}$, and conclude that $P_{1}$ is very likely a genetic variant of SARS-Cov-2. Then, if they repeat the same process with the second sequence as query $P_{2}$, they will find that only $3.12\%$ of the $q$-grams of $P_{2}$ are 0-consistent with $\mathcal{S}$. Thus, they will conclude that $P_{2}$ is unlikely to be a genetic variant of SARS-Cov-2.
Apart from practical applications, SM is motivated from a theory standpoint: SM can be reduced to RM leading to an $\mathcal{O}(N\sqrt{N})$-time baseline (see Section 3 for the details). This gives rise to two fundamental questions: (Q1) Can we solve SM significantly faster? (Q2) Can we solve problems that generalize SM to more complex graph types efficiently?
@@
    Theorem 1 answers Q1 affirmatively. The algorithm to construct the data structure in Theorem 1 computes the answer $(c_{v}^{\max},f_{v}^{\max})$, for every node $v$ of the input tree $\mathcal{T}$. It works by first splitting $\mathcal{T}$ into a forest of $\Delta$ trees $\mathcal{T}_{0},\ldots,\mathcal{T}_{\Delta-1}$, such that each tree $\mathcal{T}_{i}$, for $i\in[0,\Delta)$, has all its leaves colored $i$. Every node of each tree $\mathcal{T}_{i}$ is associated with one node of $\mathcal{T}$. Then, the algorithm makes a bottom-up traversal of $\mathcal{T}$, and, for each node $v$ of $\mathcal{T}$, it combines the color frequency information of the children of $v$ in $\mathcal{T}$ with the information coming from the nodes associated with $v$ in the trees $\mathcal{T}_{0},\ldots,\mathcal{T}_{\Delta-1}$. The efficiency of this algorithm is based on the fact that the total size of all trees is $\mathcal{O}(N)$, and on that it employs: (I) an efficient data structure for answering Lowest Common Ancestor (LCA) queries [BF00]; and (II) an efficient algorithm for tree traversal that exploits LCA information [Kas+01]. See Section 4.
@@
    We show how the above-mentioned string-processing problems, namely $1$-DR, UPM, and CQS, can be solved in optimal time using Theorem 1 via linear-time reductions to SM. We remark that via the reduction from $1$-DR to SM, we solve $1$-DR by constructing in linear time a data structure that supports queries in optimal time. The existing data structures [NN12, Hon+14, NN17, NN25] for the more general $k$-DR problem clearly work (with $k=1$) for our problem but, to the best of our knowledge, they do not admit an linear-time construction. Their focus is on obtaining theoretically good space-query time trade-offs. We also remark that the query and preprocessing time we achieve for the CQS problem are optimal. A baseline alternative approach is to construct $\Delta$ suffix trees, one for each string $S$ in the input collection $\mathcal{S}$, and to find if each $q$-gram $Q$ of the query is $\varepsilon$-consistent with $\mathcal{S}$ after matching it to each suffix tree to compute $|\textsf{occ}_{S}(Q)|$. This approach is prohibitively expensive, as it may take $\Omega(\Delta q|Q|)$ time for a query of length $|Q|$ and $\Delta$ is typically in the order of thousands. See Section 5.
@@
    We present experiments on $4$ real datasets from different domains showing that our algorithm is at least one order of magnitude faster and uses significantly less memory compared to three natural baselines. For example, it processes a dataset whose tree has over $7$ billion nodes in less than $20$ minutes, while the most time- and space-efficient baseline needs about $6.5$ hours and uses $28\%$ more memory. We also present case studies showing the usefulness of our approach in the UPM and CQS problems. In UPM, our approach discovers patterns that reveal behavioral preferences about movies, books, or products, which prevail in different user subpopulations and are reflected in the literature, and in CQS it distinguishes between DNA sequences that belong to different entities. See Section 9.

### v2

@@ Abstract
The mode of a collection of values (i.e., the most frequent value in the collection) is a key summary statistic. Finding the mode in a given range of an array of values is thus of great importance, and constructing a data structure to solve this problem is in fact the well-known Range Mode problem. In this work, we introduce the Subtree Mode (SM) problem, the analogous problem in a leaf-colored tree, where the task is to compute the most frequent color in the leaves of the subtree of a given node. SM is motivated by several applications in domains such as text analytics and biology, where the data are hierarchical and can thus be represented as a (leaf-colored) tree. Our central contribution is a time-optimal algorithm for SM that computes the answer for every node of an input $N$-node tree in $\mathcal{O}(N)$ time. We further show how our solution can be adapted for node-colored trees, or for computing the $k$ most frequent colors, for any given $k=\mathcal{O}(1)$, in the optimal $\mathcal{O}(N)$ time. Moreover, we prove that a similarly fast solution for when the input is a sink-colored directed acyclic graph instead of a leaf-colored tree is highly unlikely. Our experiments on real datasets with trees of up to $7.3$ billion nodes demonstrate that our algorithm is faster than baselines by at least one order of magnitude and much more space efficient. They also show that it is effective in pattern mining, sequence-to-database search, and biology applications.
@@ 1 Introduction
A key summary statistic of a collection of values is its mode (i.e., the most frequent value in the collection) [DBLP:books/mk/HanKP2011]. Finding the mode in a given range of values of an array (e.g., a window of a sequence) is thus of great importance. In fact, constructing a data structure to solve this problem is the well-known Range Mode (RM) problem [DBLP:journals/njc/KrizancMS05, DBLP:journals/corr/abs-1101-4068, DBLP:conf/icalp/GreveJLT10, DBLP:conf/soda/WilliamsX20]. For instance, RM allows finding the most frequently purchased item over a certain time period [sea23], the most frequent $q$-gram (i.e., length-$q$ substring) occurring in a genomic region [DBLP:journals/bioinformatics/AyadPP18], or the most frequent value of an attribute in a range of database tuples [DBLP:books/mk/HanKP2011].
@@
We start by motivating SM from a theory standpoint. SM can be easily reduced to RM and by using the state-of-the-art algorithm for RM [DBLP:journals/corr/abs-1101-4068], we obtain an $\mathcal{O}(N\sqrt{N})$-time baseline for SM (see Section 3 for the details). Unfortunately, there is strong evidence suggesting that a significantly faster algorithm for RM is rather unlikely [DBLP:journals/corr/abs-1101-4068]. This gives rise to two fundamental questions: (Q1) Can we solve SM directly and significantly faster? (Q2) Can we solve problems that generalize SM to more complex graph types efficiently?
We next motivate SM from a practical standpoint. SM is motivated by several applications in domains such as text analytics and biology. We sketch some of these below.
SNP-based Phylogenetic Tree Annotation. Single-nucleotide polymorphisms (SNPs) are genetic variations at specific nucleotide sites within a species’ genome. SNPs are linked to diseases such as cancer, Alzheimer’s disease, and other inherited disorders [Kostem2011, pnas04]. Owing to their physical proximity to disease-associated loci, SNP alleles are frequently co-inherited with pathogenic variants across generations, reflecting the principle of genetic linkage [Syvanen2001AccessingGV]. Phylogenetic trees show evolutionary relationships between organisms; a leaf represents an organism from which a DNA sequence has been obtained, and an internal node represents a common ancestor of all the organisms that correspond to its leaf descendants [phylobook]. The leaves are often annotated manually with categorical values related to SNPs (e.g., types of diseases [10.1371/journal.pone.0005022]), and each internal node with a value summarizing the values of its subtree [10193, plospathogens, pnas2025, 10.1093/nar/gkae268]. We can model such a phylogenetic tree as $\mathcal{T}$ and color each leaf of $\mathcal{T}$ according to the SNP-related value of its corresponding leaf in the phylogenetic tree. Then, the mode of $v$ identifies the most prevalent SNP-related value among the group of organisms corresponding to $v$, supporting interpretation and hypothesis generation about evolutionary processes [mbe25]. Thus, SM can reveal significant associations, such as clade‑specific markers [Coll2014] and genotype–phenotype links [10.1093/bioinformatics/btad215, Arora2017OriginSyphilis], between subtrees and SNP-related values. This is valuable in studies of pathogen spread [plospathogens] and antibiotic resistance evolution [Arora2017OriginSyphilis]. For example, coloring the phylogenetic tree in [10193, Fig. 3b] by “continent of origin” and solving SM identifies subtrees where Ethiopian samples are prevalent, which is central in the study of [plospathogens]. Similarly, coloring the phylogenetic tree in [10193, Fig. S1] by SNP state at selected genomic positions (respectively, by antibiotic resistance status) identifies subtrees where particular SNP states (respectively, resistant samples) predominate in the data of the study of [Arora2017OriginSyphilis], which is crucial for detecting genotype-phenotype links. There are widely-used systems for annotating phylogenetic tree nodes with values (e.g., [itol6, 10193, icytree]) but one must manually compute the most frequent colors (i.e., solve SM) and perform the annotation in these systems, whereas our approach is automatic and efficient.
@@
Fig. 1 shows an SNP-based phylogenetic tree alike the ones used to study the Influenza A virus in [10.1371/journal.pone.0005022]. The leaves are annotated with Influenza A subtypes (H1N1, H1N2, H5N1, and H5N8), and the internal nodes represent lineages (subtrees sharing specific patterns of SNPs). An internal node with .1 and .2 corresponds to a subtype in the Eastern and Western Hemisphere, respectively. We model the tree in Fig. 1 as $\mathcal{T}$ in our SM problem and assign colors 0 (blue), $1$ (green), $2$ (red), and $3$ (orange) to the leaves of subtype H1N1, H1N2, H5N1, and H5N8, respectively. Thus, the number of nodes $N$ of $\mathcal{T}$ is $20$, and the number of colors $\Delta$ is $4$. Given a query consisting of the node H1, SM outputs 0, as the subtree rooted at this node has more blue leaves than green.
Top-11 Document Retrieval ($1$-DR). In the $1$-DR problem [DBLP:conf/soda/NavarroN12, DBLP:journals/jacm/HonSTV14, DBLP:journals/siamcomp/NavarroN17, DBLP:conf/soda/0001N25], a collection $\mathcal{S}$ of $\Delta$ documents (strings), $S_{0},\ldots,S_{\Delta-1}$, is given for preprocessing, and we are asked to answer queries of the following type: given a query pattern $P$, output the string in $\mathcal{S}$ in which $P$ occurs most frequently as a substring. $1$-DR can be reduced to SM. We preprocess $\mathcal{S}$ by first constructing its suffix tree (i.e., the compacted trie of the suffixes of the string $S_{0}\mathdollar_{0}\ldots S_{\Delta-1}\mathdollar_{\Delta-1}$, where $\mathdollar_{i}$, for each $i\in[0,\Delta)$, is a unique delimiter) [DBLP:conf/focs/Weiner73] and then coloring the leaves corresponding to the suffixes starting in $S_{i}\mathdollar_{i}$ with color $i$. This is the leaf-colored tree in the constructed instance of SM. For a query pattern $P$, we spell $P$ on the suffix tree, arriving at a node $v$, compute the mode of $v$, say $i$, and output $S_{i}$ as the answer. The example below illustrates how we solve $1$-DR via SM.
@@
Uniform Pattern Mining (UPM). Consider two strings, one comprised of male-targeted ads and another comprised of female-targeted ads, and that we mine a pattern “strong leader” which occurs much more frequently in the former string. Ads with this pattern may perpetuate gender stereotypes and influence how opportunities or messages are presented to different ad viewer groups, causing discrimination and bias in decision-making or perception. To prevent this, in the spirit of statistical parity [parity] [^1] This fairness measure requires the probability distributions of outcomes to be similar across all subpopulations of a population., we can mine patterns with “similar” frequencies in all strings of an input collection by formulating and solving the following problem, called Uniform Pattern Mining (UPM): Given a collection $\mathcal{S}$ of $\Delta$ strings, $S_{0},\ldots,S_{\Delta-1}$, and an integer $\varepsilon\geq 0$ specified based on domain knowledge, UPM asks for all strings (patterns) whose frequencies in any pair of strings in $\mathcal{S}$ differ by at most $\varepsilon$. When the strings in $\mathcal{S}$ represent different subpopulations of a user population and $\varepsilon$ is “small”, such patterns prevent the discrimination of these subpopulations. On the other hand, when $\varepsilon$ is “large”, those patterns with large differences in their frequencies reveal behavioral preferences that prevail in user subpopulations (e.g., they may represent movie genres viewed by much more men than women). The UPM problem is solved via a reduction to SM.
@@
Consistent Query String (CQS). In the CQS problem, a collection $\mathcal{S}$ of $\Delta$ strings, $S_{0},\ldots,S_{\Delta-1}$, is given for preprocessing, and we are asked to quantify how similar a query string $P$ is to the strings in $\mathcal{S}$. This can be achieved by counting the number of distinct $q$-grams $Q$ of $P$ whose frequency in $P$ is in the interval ${[\min_{S\in\mathcal{S}}|\textsf{occ}_{S}(Q)|-\varepsilon,}\max_{S\in\mathcal{S}}|\textsf{occ}_{S}(Q)|+\varepsilon]$, where $|\textsf{occ}_{S}(Q)|$ is the frequency of $Q$ in string $S$ and $\varepsilon$ is a user-specified parameter capturing approximate consistency. We call each such $q$-gram $\varepsilon$-consistent with $\mathcal{S}$. Clearly, $P$ is similar to the strings in $\mathcal{S}$ if most of its $q$-grams are $\varepsilon$-consistent with $\mathcal{S}$. The CQS problem is particularly relevant for databases of highly-similar strings, which are common in genomics, as they are constructed over collections with a shared evolutionary history or a common function [shuy2017gisaid, DBLP:journals/nar/MistryCWQSSTPRR21]. For example, the CQS problem can be used for identifying viral species, a core challenge in virology [vir1, vir2, vir3, vir4, vir5]. State-of-the-art methods for this task [vir3, vir4, vir5] perform $q$-gram based sequence classification but need separate training for each desired $q$ and decide class membership for a pattern (e.g., SARS-Cov-2 vs. Influenza). On the contrary, a data structure for CQS is built once and queried for any $q$, and it provides numerical scores, which is more informative when sequences have errors, as is often the case in practice.
@@
Consider two DNA sequences, one that is a genetic variant of SARS-Cov-2 [ncbi_sars_cov_2_2024] and another that is a genetic variant of the Influenza A virus [ncbi_influenza_a_2025]. Suppose that a biologist does not know whether each of these sequences is a genetic variant of SARS-CoV-2 and wants to check this. They can use the first sequence as query $P_{1}$ in a database $\mathcal{S}$ comprised of $2,000$ different genetic variants of the SARS CoV-2 virus [ncbi_sars_cov_2_2024], and solve CQS for $\varepsilon=0$ and $q=4$. They will find that $99.9\%$ of the $q$-grams of $P_{1}$ are 0-consistent with $\mathcal{S}$, and conclude that $P_{1}$ is very likely a genetic variant of SARS-Cov-2. Then, if they repeat the same process with the second sequence as query $P_{2}$, they will find that only $3.12\%$ of the $q$-grams of $P_{2}$ are 0-consistent with $\mathcal{S}$. Thus, they will conclude that $P_{2}$ is unlikely to be a genetic variant of SARS-Cov-2.
@@
    Theorem 1 answers Q1 affirmatively. The algorithm to construct the data structure in Theorem 1 computes the answer $(c_{v}^{\max},f_{v}^{\max})$, for every node $v$ of the input tree $\mathcal{T}$. It works by first splitting $\mathcal{T}$ into a forest of $\Delta$ trees $\mathcal{T}_{0},\ldots,\mathcal{T}_{\Delta-1}$, such that each tree $\mathcal{T}_{i}$, for $i\in[0,\Delta)$, has all its leaves colored $i$. Every node of each tree $\mathcal{T}_{i}$ is associated with one node of $\mathcal{T}$. Then, the algorithm makes a bottom-up traversal of $\mathcal{T}$, and, for each node $v$ of $\mathcal{T}$, it combines the color frequency information of the children of $v$ in $\mathcal{T}$ with the information coming from the nodes associated with $v$ in the trees $\mathcal{T}_{0},\ldots,\mathcal{T}_{\Delta-1}$. The efficiency of this algorithm is based on the fact that the total size of all trees is $\mathcal{O}(N)$, and on that it employs: (I) an efficient data structure for answering Lowest Common Ancestor (LCA) queries [DBLP:conf/latin/BenderF00]; and (II) an efficient algorithm for tree traversal that exploits LCA information [DBLP:conf/cpm/KasaiLAAP01]. See Section 4.
@@
    We show how the above-mentioned string-processing problems, namely $1$-DR, UPM, and CQS, can be solved in optimal time using Theorem 1 via linear-time reductions to SM. We remark that via the reduction from $1$-DR to SM, we solve $1$-DR by constructing in linear time a data structure that supports queries in optimal time. The existing data structures [DBLP:conf/soda/NavarroN12, DBLP:journals/jacm/HonSTV14, DBLP:journals/siamcomp/NavarroN17, DBLP:conf/soda/0001N25] for the more general $k$-DR problem clearly work (with $k=1$) for our problem but, to the best of our knowledge, they do not admit an linear-time construction. Their focus is on obtaining theoretically good space-query time trade-offs. We also remark that the query and preprocessing time we achieve for the CQS problem are optimal. A baseline alternative approach is to construct $\Delta$ suffix trees, one for each string $S$ in the input collection $\mathcal{S}$, and to find if each $q$-gram $Q$ of the query is $\varepsilon$-consistent with $\mathcal{S}$ after matching it to each suffix tree to compute $|\textsf{occ}_{S}(Q)|$. This approach is prohibitively expensive, as it may take $\Omega(\Delta q|Q|)$ time for a query of length $|Q|$ and $\Delta$ is typically in the order of thousands. See Section 5.
@@
    We present experiments on $4$ real datasets from different domains showing that our algorithm is at least one order of magnitude faster and uses significantly less memory compared to three natural baselines. For example, it processes a dataset whose tree has over $7$ billion nodes in less than $20$ minutes, while the most time- and space-efficient baseline needs about $6.5$ hours and uses $28\%$ more memory. We also show the efficiency of our approach in $k$-DR and its usefulness in UPM, CQS, and a phylogenetic tree annotation application. In UPM, our approach discovers patterns that reveal behavioral preferences about movies, books, or products, which prevail in different user subpopulations and are reflected in the literature, in CQS it efficiently distinguishes between DNA sequences that belong to different entities, and in the last application it helps genotype-phenotype links detection. See Section 9.